OpenStudy (anonymous):

What is the LCD of 2/x + x/x+1 + X+1/3x?

OpenStudy (anonymous):

\[\frac{ 2 }{ x } + \frac{ x }{ x+1 } + \frac{ x+1 }{ 3x }\]